# LoRA-Edit: Controllable First-Frame-Guided Video Editing via Mask-Aware LoRA Fine-Tuning

### v3

@@ Abstract
Video editing using diffusion models has achieved remarkable results in generating high-quality edits for videos. However, current methods often rely on large-scale pretraining, limiting flexibility for specific edits. First-frame-guided editing provides control over the first frame, but lacks flexibility over subsequent frames. To address this, we propose a mask-based LoRA (Low-Rank Adaptation) tuning method that adapts pretrained Image-to-Video (I2V) models for flexible video editing. Our approach preserves background regions while enabling controllable edits propagation. This solution offers efficient and adaptable video editing without altering the model architecture. To better steer this process, we incorporate additional references, such as alternate viewpoints or representative scene states, which serve as visual anchors for how content should unfold. We address the control challenge using a mask-driven LoRA tuning strategy that adapts a pre-trained image-to-video model to the editing context. The model must learn from two distinct sources: the input video provides spatial structure and motion cues, while reference images offer appearance guidance. A spatial mask enables region-specific learning by dynamically modulating what the model attends to, ensuring that each area draws from the appropriate source. Experimental results show our method achieves superior video editing performance compared to state-of-the-art methods.
@@ 1 Introduction
Recent advances of diffusion models (Rombach et al., 2022; Lipman et al., 2023) has demonstrated unprecedented improvement in high-quality video generation (Yang et al., 2025b; Kong et al., 2024; Wang et al., 2025; HaCohen et al., 2024). Based on foundation video generation model, video editing has also experienced dramatic improvement (Jiang et al., 2025; Hu et al., 2025), widely used in creative, commercial, and scientific uses these days. Still, these video editing models often require a heavy-compute finetuning, with a large set of training data. This makes them very expensive to extend a new editing type, and less flexible to new applications.
@@
In contrast, first-frame-guided video editing (Ouyang et al., 2024; Ku et al., 2024) offers a promising path toward flexible video manipulation. In this paradigm, users edit the first frame arbitrarily, either using image AI tools or traditional editing software. These edits are then propagated to the entire sequence, enabling flexible video manipulation without being constrained by dataset-specific training.
@@
While first-frame-guided solutions allows flexible editing, it provides limited control of remaining frames. For instance, given a video of a blooming flower, the user can edit the flower in the first frame, but cannot control how the flower blooms in the following frames. Similarly, when an object rotates to a novel viewpoint, the user cannot control appearance of the disoccluded region of this object, as shown in the bottom row of Fig. 1, where the handbag rotates to a new angle. In addition, the first frame edits may diffuse into unedited regions, resulting undesirable background changes. The inability to control later frames limits editing flexibility and necessitates methods that not only retain the flexibility of first-frame-guided editing, but also support control throughout the video.
@@
A simple solution is per-video finetuning of a pre-trained image-to-video (I2V) model (Kong et al., 2024; Wang et al., 2025). By finetuning the model using LoRA (Hu et al., 2022) on an input video, the model will learn content motion, which can be applied to an edited first frame. This allows the edit to propagate in a temporally consistent way. However, this naive approach lacks finer control—it cannot distinguish between regions that should change and those that should stay, nor does it ensure that the appearance of the edited region remains controllable as it moves and deforms over time.
@@
In this work, we build a flexible video editing model by expanding this naive edit propagation approach with an additional mask, which controls which regions of the video remain unchanged and which are modified. Recent I2V models (Kong et al., 2024; Wang et al., 2025) are designed to generate videos from a single image, but they can also process video sequences, with a built-in masking mechanism to control which regions are preserved or modified during inference. Typically, this mask preserves the first frame while generating the subsequent frames. However, we further observe that the mask has greater potential for more precise control over video content. To leverage this, we apply LoRA to fine-tune the model on the input video with the edited region masked. This allows the model to learn how to use the mask for controlling local content generation while preserving the background. After LoRA training, the model can effectively apply the mask, ensuring that unedited areas remain unchanged during inference.
More importantly, the mask-based control enables LoRA to learn flexible information from the training data, adapting to various editing conditions. For example, when the mask preserves the first frame and generates the subsequent frames, LoRA learns motion patterns from the input video, allowing for temporally consistent edits. On the other hand, with a fully blank mask, LoRA can focus on learning the appearance of specific objects from a single frame, providing detailed control over visual content. Our approach offers a simple and effective solution for video editing by leveraging LoRA’s capabilities, without modifying the model architecture, and maintaining high flexibility through the combination of different conditions. Extensive experiments demonstrate that our method achieves superior performance over previous state-of-the-art approaches in both qualitative and quantitative evaluations. Please visit our project page for video editing examples.
@@ 2 Related Works
Video Editing with Diffusion Models
The success of video diffusion models has spurred extensive research into video editing. Early works adapt the image diffusion network and training paradigm to video generation and editing. Tune-A-Video Wu et al. (2023) explores the concept of one-shot tuning in video editing. Fairy Wu et al. (2024) edits keyframes utilizing a 3D spatio-temporal self-attention extended from a T2I diffusion model. VidToMe Li et al. (2024) introduces image editing approaches (e.g., ControlNet Zhang et al. (2023)) to video generation. Animatediff Guo et al. (2023) decouples the appearance and motion learning during video editing. SAVE Song et al. (2024) chooses to fine-tune the feature embeddings that directly reflect semantic information. Another line of work manipulates the hidden features to edit a video. Video-P2P Liu et al. (2024) and Vid2Vid-Zero Wang et al. (2023) employ cross-attention map injection and null-text inversion for video editing. TokenFlow Geyer et al. (2023) leverages motion-based feature injection, and FLATTEN Cong et al. (2023) further introduces optical flow for better injection. Other methods Chen et al. (2023); Yang et al. (2023) explore latent initialization and latent transition in video diffusion models. Dragvideo Deng et al. (2024) achieves interactive drag-style video editing by introducing point conditioning. Recently, VACE Jiang et al. (2025) has shown promising video editing ability by large-scale conditional video diffusion training. Although large video editing diffusion models achieve impressive results, they often struggle with inaccurate identity preservation and suboptimal performance on out-of-domain test cases. In contrast, our method effectively leverages powerful video priors while efficiently learning content from both the reference image and the source video.
First-Frame Guided Video Editing
First-frame guided editing has emerged as a mainstream video editing approach, with AnyV2V Ku et al. (2024) and I2VEdit Ouyang et al. (2024) as representative methods. These approaches decompose video editing into two stages: (i) editing the first frame using existing image methods, and (ii) propagating edits to remaining frames using motion-conditioned image-to-video diffusion models. AnyV2V reconstructs motion via DDIM sampling, injecting temporal attention and spatial features from the original video. I2VEdit enhances this by learning coarse motion through per-clip LoRA and refining appearance using attention difference masks. While this decoupled framework benefits from advances in both image editing and video generation, the lack of explicit constraints often leads to diluted edits during propagation, manifesting as foreground inconsistencies and background leakages.
@@ 3.1 LoRA’s First Step: A Simple Solution for Video Editing
To achieve this basic objective, we insert LoRA (Low-Rank Adaptation) (Hu et al., 2022) modules into the self-attention and cross-attention layers of the I2V model(Wang et al., 2025) and optimize them on the input video to capture its motion pattern. During training, the model is conditioned on the original first frame and a textual prompt composed of a fixed special token concatenated with the caption generated for using Florence-2 (Xiao et al., 2024) (i.e., + ). The model is supervised to reconstruct the full input video . Following the denoising objective of the I2V diffusion model(Wang et al., 2025), we optimize the LoRA adapters by minimizing the error between the network noise prediction and the Gaussian noise injected into the latent video:
where is the sampled noise, and denotes the noise prediction network with LoRA parameters . is the VAE encoder that maps the target video input to the latent space.
@@ 3.2 The Mask’s Hidden Power: Exploring I2V Model Capabilities
Although naive edit propagation ensures motion coherence, it lacks control over the content of subsequent frames. To address this, we leverage the conditioning mechanisms in recent I2V models (Wang et al., 2025; Kong et al., 2024). To introduce the first frame as the guidance for video generation, these models incorporate two additional conditions for the denoising network: a pseudo-video and a binary spatiotemporal mask . The pseudo-video is constructed by concatenating the first frame with zero-placeholder frames. The binary mask is designed so that 1 indicates the preserved frame and 0 represents the frames to be generated, with the first frame set to 1 and all subsequent frames set to 0.
This paradigm can be extended to video-to-video generation by replacing the pseudo-video condition with actual video frames, enabling the model to accept an entire video sequence as input. In this setting, the binary spatiotemporal mask , originally designed to preserve only the first frame, can now be repurposed as a more flexible mechanism that selectively controls which regions are retained and which are regenerated across space and time. To assess the generalization capabilities of the masking mechanism, we evaluate several binary mask configurations. In each case, the mask is applied to the input video to construct , where regions marked as zero are regenerated and the rest are preserved.
Default I2V Configuration. In this configuration, the I2V model preserves the first frame and generates the remaining frames based on the first frame. As shown in the first row of Fig. 2, this default mask setup leads the model to synthesize motion across the entire sequence.
Case #1: No Preservation. In this case, the model is configured with an all-zeros mask across all frames, meaning none of the original video content is preserved. As shown in the second row of Fig. 2, this setup can force the model to generate the appearance for the entire video.
Case #2: All Preservation. In this case, the mask preserves the content from the input video while setting the generated frames to zero, ensuring that the original video sequence is maintained in the preserved regions. As shown in the third row of Fig. 2, although the overall structure of the original video content is effectively preserved, artifacts appear in areas with discontinuous motion. Inspired by this, further LoRA fine-tuning could potentially improve its ability to preserve input video.
Case #3: Selective Preservation. In this case, a spatially varying mask is used to preserve the background while allowing the foreground to be generated. As shown in the fourth row of Fig. 2, the model struggles to synthesize coherent foreground content. This suggests that the pretrained I2V model lacks inpainting capabilities, likely due to limited exposure to spatially varying masks during training. More importantly, the model has difficulty handling the interaction between edits and the background, which we aim to improve by fine-tuning on the input video using LoRA.
@@ 3.3 Unlocking Editing Flexibility: Mask-Guided LoRA
Building on this exploration, we modify the spatiotemporal mask to enable more flexible video edits. Combined with LoRA fine-tuning, the mask serves two complementary roles. First, it improves the I2V model’s alignment with mask constraints, allowing flexible control over which regions are edited or preserved. Second, it acts as a signal guiding LoRA to learn specific patterns from the training data, such as motion from video sequences or appearance from images. Specifically, we modify the training loss in Eq. 3.1 to introduce the conditioning video and mask:
As shown in Fig. 3, by configuring , , and in different ways, we enable flexible video editing through LoRA, which will be described in detail in the following sections.
3.3.1 Disentangling Edits and Background.
Many first-frame edits alter only part of the frame, yet they impose two contrasting demands: the edited region must keep evolving, while the untouched background should stay locked. When both demands run through the same generative pathway, they collide—the drive to preserve the background can freeze the edit, while the drive to propagate the edit can leak changes into the background.
To effectively disentangle edited regions from the background, we carefully adjust the spatiotemporal mask and the conditioning video during LoRA fine-tuning. The mask is set to ones for the first frame to preserve it as the reference, and for subsequent frames, is adjusted to mark unedited regions with ones (to be preserved) and edited regions with zeros (to be generated). The pseudo-video is created by applying the mask to the input video, setting the regions marked as zero in to be empty, while leaving the rest unchanged. The objective is set to the input video during LoRA fine-tuning. This configuration allows the model to focus on generating the edited content while locking the unedited regions. At inference time, when editing the first frame (replacing with ), we use the same as during LoRA training, while has its first frame replaced by the edited version .
While the pre-trained I2V model struggles with selective preservation, LoRA on a single video alone surprisingly enables the model to learn effective mask-guided inpainting priors. We speculate this is due to the diffusion transformer processing inputs as discretized tokens, with a spatially varying mask sharing a similar token-level representation, making the adaptation not very difficult.
3.3.2 Appearance Control in Propagated Edits.
@@
During LoRA fine-tuning, we use an edited frame as the target . The conditioning input is constructed using the pre-edited frame by masking out the edited regions. The mask marks the preserved background areas with ones and the edited regions with zeros. If multiple frames are involved, each frame is treated as a separate sample to avoid including motion information. This configuration allows the model to learn how edited content should appear in context, guided by both the surrounding background and the user-provided modification.
Unlike methods that directly feed edited frames as inputs during inference (Yang et al., 2025a; Jamriska, 2018), we do not require the edited frame to remain exactly the same during inference. Instead, the edited frame is used only during training to guide how edits should appear. At inference time, the model generates content based on learned patterns and context, allowing it to adapt edits smoothly across frames, even when the edits do not adhere to strict temporal alignment.

### v6

# Controllable First-Frame-Guided Video Editing via Mask-Aware LoRA Fine-Tuning

@@ Abstract
Video editing using diffusion models has achieved remarkable results in generating high-quality edits for videos. However, current methods often rely on large-scale pretraining, limiting flexibility for specific edits. First-frame-guided editing provides control over the first frame, but lacks fine-grained control over the edit’s subsequent temporal evolution. To address this, we propose a mask-based LoRA (Low-Rank Adaptation) tuning method that adapts pretrained Image-to-Video models for flexible video editing. Our key innovation is using a spatiotemporal mask to strategically guide the LoRA fine-tuning process. This teaches the model two distinct skills: first, to interpret the mask as a command to either preserve content from the source video or generate new content in designated regions. Second, for these generated regions, LoRA learns to synthesize either temporally consistent motion inherited from the video or novel appearances guided by user-provided reference frames. This dual-capability LoRA grants users control over the edit’s entire temporal evolution, allowing complex transformations like an object rotating or a flower blooming. Experimental results show our method achieves superior video editing performance compared to baseline methods. The code and video results are available at our project website: https://cjeen.github.io/LoRAEdit.
@@ 1 Introduction
Recent advances in diffusion models (Rombach et al., 2022; Lipman et al., 2023) have demonstrated unprecedented improvement in high-quality video generation (Yang et al., 2025b; Kong et al., 2024; Wang et al., 2025; HaCohen et al., 2024). Based on foundational video generation models, video editing has experienced dramatic improvement (Jiang et al., 2025; Hu et al., 2025), now finding wide application in scientific and creative fields (Yang et al., 2024). Still, these video editing models often require computationally intensive finetuning, with a large set of training data. This makes them expensive to extend to a new editing type, and less flexible for new applications.
@@
In contrast, first-frame-guided video editing (Ouyang et al., 2024; Ku et al., 2024) offers a promising path toward flexible video manipulation. In this paradigm, users can edit the first frame arbitrarily, either using image AI tools or traditional editing software. These edits are then propagated to the entire sequence, enabling flexible video manipulation without being constrained by dataset-specific training.
@@
While first-frame-guided solutions allow flexible editing, they only provide limited control of remaining frames. For instance, given a video of a blooming flower, the user can edit the flower in the first frame, but cannot control how the flower blooms in the following frames. Similarly, when an object rotates to a novel viewpoint, the user cannot control the disoccluded region. In addition, the first frame edits may diffuse into unedited regions, resulting in undesirable background changes. The inability to control later frames limits editing flexibility and necessitates methods that not only retain the flexibility of first-frame-guided editing, but support control throughout the video.
@@
A simple solution is per-video finetuning of a pre-trained image-to-video (I2V) model (Kong et al., 2024; Wang et al., 2025). By finetuning the model using LoRA (Hu et al., 2022) on a source video, the model will learn content motion. This allows the edit to propagate in a temporally consistent way. However, this naive approach lacks finer control—it cannot distinguish between regions that should change and those that should stay, nor does it ensure the edited region’s appearance remains controllable as it moves and deforms over time, requiring the synthesis of unseen appearance.
@@
In this work, we build a flexible video editing model by expanding this naive edit propagation approach with an additional mask, which controls which regions of the video remain unchanged and which are modified. Recent I2V models (Kong et al., 2024; Wang et al., 2025) are designed to generate videos from a single image, but they can also process video sequences, with a built-in masking mechanism to control which regions are preserved or modified during inference. Typically, this is only a temporal mask, which preserves the first frame while generating the subsequent frames. However, we observe the mask has greater potential for more precise spatial control over video content. To leverage this, we apply LoRA to fine-tune the model on the input video with the edited region masked. This allows the model to learn how to interpret a flexible spatiotemporal mask as a command to preserve content from the source video or generate new content in designated regions. After LoRA training, the model can effectively apply the mask, leaving unedited areas unchanged.
More importantly, our mask-guided control empowers LoRA to learn selectively, adapting to the specific demands of the edit. This flexibility is illustrated by the blooming flower example in Fig. 1. First, by configuring the mask for motion learning, LoRA learns the flower’s blooming motion from the source video (second row). Then, to control the flower’s appearance as it blooms, the mask is reconfigured for appearance learning, directing LoRA to capture the target appearance from an additional reference, such as the bloomed, vibrant red rose (third row). This dual capability allows our method to synthesize a controllable transformation, creating a video where the flower not only moves correctly but evolves into the desired state, a feat unattainable with naive first-frame guidance.
Our approach offers a simple and effective solution for video editing by leveraging LoRA’s capabilities, without modifying the model architecture, and maintaining high flexibility through the combination of different conditions. Experimental results demonstrate that our method achieves superior performance over state-of-the-art approaches in both qualitative and quantitative evaluations.
@@ 2 Related Works
Video Editing with Diffusion Models. The success of video diffusion models has spurred extensive research into video editing. Early works adapt the image diffusion network and training paradigm to video generation and editing. Tune-A-Video Wu et al. (2023) explores the concept of one-shot tuning in video editing. Fairy Wu et al. (2024) edits keyframes utilizing a 3D spatio-temporal self-attention extended from a T2I diffusion model. VidToMe Li et al. (2024) introduces image editing approaches (e.g., ControlNet Zhang et al. (2023)) to video generation. Animatediff Guo et al. (2023) decouples the appearance and motion learning during video editing. SAVE Song et al. (2024) chooses to fine-tune the feature embeddings that directly reflect semantic information. Another line of work manipulates the hidden features to edit a video. Video-P2P Liu et al. (2024) and Vid2Vid-Zero Wang et al. (2023) employ cross-attention map injection and null-text inversion for video editing. TokenFlow Geyer et al. (2023) leverages motion-based feature injection, and FLATTEN Cong et al. (2023) further introduces optical flow for better injection. Other methods Chen et al. (2023); Yang et al. (2023) explore latent initialization and latent transition in video diffusion models. Dragvideo Deng et al. (2024) achieves interactive drag-style video editing by introducing point conditioning. Recently, VACE Jiang et al. (2025) has shown promising video editing ability by large-scale conditional video diffusion training. Although large video editing diffusion models achieve impressive results, they often struggle with inaccurate identity preservation and suboptimal performance on out-of-domain test cases. In contrast, our method effectively leverages powerful video priors while efficiently learning content from both the reference image and the source video.
First-Frame Guided Video Editing. First-frame guided editing has emerged as a mainstream video editing approach, with AnyV2V Ku et al. (2024) and I2VEdit Ouyang et al. (2024) as representative methods. These approaches decompose video editing into two stages: (i) editing the first frame using existing image methods, and (ii) propagating edits to remaining frames using motion-conditioned image-to-video diffusion models. AnyV2V reconstructs motion via DDIM sampling, injecting temporal attention and spatial features from the original video. I2VEdit enhances this by learning coarse motion through per-clip LoRA and refining appearance using attention difference masks. While this decoupled framework benefits from advances in both image editing and video generation, the lack of explicit constraints often leads to diluted edits during propagation, manifesting as foreground inconsistencies and background leakages.
@@ 3.1 LoRA’s First Step: A Simple Solution for Video Editing
To achieve this basic objective, we insert LoRA (Hu et al., 2022) modules into the self-attention and cross-attention layers of the I2V model(Wang et al., 2025) and optimize them on the input video to capture its motion pattern. During training, the model is conditioned on the original first frame and a textual prompt composed of a fixed special token concatenated with the caption generated for using Florence-2 (Xiao et al., 2024) (i.e., + ). The model is supervised to reconstruct the full input video . Following the flow matching objective of the I2V model (Wang et al., 2025), we optimize the LoRA adapters by minimizing the error between the network velocity prediction and the target flow drift:
where represents the sampled Gaussian noise, and denotes the latent representation of the source video encoded by the VAE . Here, is the velocity prediction network with LoRA parameters , and is the time step. The term serves as the flow target, guiding the interpolation from data to noise .
@@ 3.2 The Mask’s Hidden Power: Exploring I2V Model Capabilities
The naive edit propagation introduced above only ensures motion coherence and lacks control over the content of subsequent frames. To address this, we leverage the conditioning mechanisms in recent I2V models (Wang et al., 2025; Kong et al., 2024). Specifically, to introduce the first frame as the guidance for video generation, these models incorporate two additional conditions for the denoising network: a pseudo-video and a binary spatiotemporal mask . The pseudo-video is constructed by concatenating the first frame with zero-placeholder frames. The binary mask is designed so that 1 indicates the preserved frame and 0 represents the frames to be generated, with the first frame set to 1 and all subsequent frames set to 0.
We extend this paradigm to video-to-video generation by replacing the pseudo-video condition with actual video frames, enabling the model to accept an entire video sequence as input. In this setting, the binary spatiotemporal mask , originally designed to preserve only the first frame, can now be repurposed as a more flexible mechanism that selectively controls which regions are retained and which are regenerated across space and time.
To assess the generalization capabilities of the masking mechanism, we evaluate several binary mask configurations, as shown in Fig. 2. In each case, the mask is applied to the input video to construct , where regions marked as zero are regenerated and the rest are preserved. The default I2V configuration preserves only the first frame and leads the model to synthesize motion across the entire sequence (first row). Exploring the extremes, an all-zeros mask that preserves none of the original content forces the model to generate the appearance for the entire video (second row). Conversely, an all-ones mask aimed at preserving all content effectively maintains the video’s overall structure but introduces artifacts in areas with discontinuous motion (third row). Finally, using a spatially varying mask to preserve the background while generating the foreground reveals a key challenge, as the model struggles to synthesize coherent foreground content (fourth row).
Analysis and Motivation. The preceding cases show the raw I2V model can handle simple, full-frame instructions but fails at nuanced, selective editing. Our key insight is to repurpose this mechanism, enabling the spatiotemporal mask to serve a dual purpose during LoRA tuning. First, we can use LoRA to reinforce the model’s response to the mask, improving its ability to execute the preservation and generation commands defined by the mask. More critically, we can use the mask to direct what LoRA learns. By masking different content during training, we can guide LoRA to focus on either the video’s underlying motion or a reference’s target appearance. This interplay between LoRA and the mask is the cornerstone of our method, detailed in the following section.
@@ 3.3 Unlocking Editing Flexibility: Mask-Guided LoRA
Building on this exploration, we modify the spatiotemporal mask to enable more flexible video edits. Combined with LoRA fine-tuning, the mask serves two complementary roles. First, it improves the I2V model’s alignment with mask constraints, allowing users to limited editing to a specific region, achieving more flexible control. Second, it also acts as a signal guiding LoRA to learn specific patterns from the training input, such as motion pattern from video sequences or appearance of an object from images. Specifically, we adapt the flow matching loss in Eq. 1 to incorporate the conditioning video and mask:
where represents the latent of the target video, and denotes the sampled noise. As shown in Fig. 3, by configuring , , and in different ways, we enable flexible video editing through LoRA, detailed in the following sections.
3.3.1 Disentangling Edits and Background
Many first-frame edits alter only a part of the frame(Ku et al., 2024; Ouyang et al., 2024), creating a conflict between two demands: the edited region must evolve, while the background must remain static. When a single generative pathway handles both, they may collide. Preserving the background can stall the edit, while propagating the edit can cause unintended background changes.
To achieve separate on control on edited regions and non-edit background, we carefully adjust the spatiotemporal mask and the conditioning video during LoRA fine-tuning. The mask is set to ones for the first frame to preserve it as the reference, and for subsequent frames, is adjusted to mark unedited regions with ones (to be preserved) and edited regions with zeros (to be generated). The pseudo-video is created by applying the mask to the input video, setting the regions marked as zero in to be empty, while leaving the rest unchanged. The objective is set to the input video during LoRA fine-tuning. This configuration allows the model to focus on generating the edited content while locking the unedited regions. At inference time, when editing the first frame (replacing with ), we use the same as during LoRA training, while has its first frame replaced by the edited version .
One interesting observation is that while a pre-trained I2V model struggles with selective editing, LoRA training on a single video alone learns effective mask-guided inpainting priors. We speculate this is due to the diffusion transformer processing inputs as discretized tokens, with a spatially varying mask sharing a similar token-level representation, making the adaptation straightfoward.
3.3.2 Appearance Control in Propagated Edits
@@
During LoRA fine-tuning, we use an edited frame as the target . The conditioning input is constructed using the pre-edited frame by masking out the edited regions. The mask marks the preserved background areas with ones and the edited regions with zeros. If multiple edited frames are used, we decouple appearance from motion by training on each frame as an isolated, static image, preventing the model from inferring false temporal dynamics between them. This configuration allows the model to learn how edited content should appear in context, guided by both the surrounding background and the user-provided modification.
Unlike methods that directly feed edited frames as inputs during inference (Yang et al., 2025a; Jamriska, 2018), we do not require the edited frame to remain the same during inference. Instead, the edited frame is used only during training to guide how edits should appear. At inference time, the model generates content based on learned patterns and context, allowing it to adapt edits smoothly across frames, even when the edits do not adhere to strict temporal alignment.